OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y less than or greater to -x - 4 Choices: (1, 10) (-1, 10) (10, 1) (1, -10)

Nnesha (nnesha):

you can graph both equation and can find solution where both lines intersect that point is your solution

Nnesha (nnesha):

https://www.desmos.com/calculator use this to graph and let me know if you wanna do other method

OpenStudy (danjs):

Take each point (x,y) and put in the numbers, which one makes both equations true?

OpenStudy (danjs):

For example (1,10)\[10 \le 1 - 5\] \[10 \le -1 - 4\] Those are false, so that point is not in the solution set for those 2 inequalities. I am assuming you mean 'less than or equal too' in the problem.

OpenStudy (anonymous):

Yes that is what I meant

OpenStudy (danjs):

Try the second point out (x,y) = (-1,10)

OpenStudy (danjs):

From the first inequality , using this point you can see this one is also false 10 is less than or equal to -1-5 nope

Nnesha (nnesha):

i guess there is a typo less than or equal to** less than or equal** to

OpenStudy (danjs):

yes, that is what it should be , he said above

OpenStudy (danjs):

or she

OpenStudy (danjs):

What do you get using the third point?

OpenStudy (anonymous):

I am confused though on how to do it

OpenStudy (danjs):

\[y \le x - 5\] third point is (10 , 1), in there x = 10, y = 1 Put that into the above inequality , and see if it is true or false

OpenStudy (danjs):

\[1 \le 10 - 5 ~~~;~~1 \le 5\] That is true, now try it in the second inequality

OpenStudy (danjs):

\[y \le -x - 4 ~~~~;~~~~1 \le -10-4 \] that is false, so the point (10,1) is not in the solution set, the point has to satisfy both of the inequalities to be part of the solution set

OpenStudy (danjs):

u see what i am doing, just putting the x and y values into the inequality, and seeing if it is true or false? last one says, 1 is less than or equal to negative 14, that is false

OpenStudy (danjs):

that leaves just the 4h point to test, and since the others were all false, the last point in the list should be true and be part of the solution set

OpenStudy (triciaal):

another approach is to set the y equal and solve for x

OpenStudy (triciaal):

also correct the question before checking the inequality